OpenStudy (anonymous):

The perimeter of a rectangle is 68 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 9 : 8. RATIOS PROBLEM please help me

OpenStudy (anonymous):

lenght = (9/(9+8))*(68/2) = 18 Breadth = (L+ B) - L = 34 - 18 =16 note that given 2*(L + B) = 68 (perimeter)